# AuxDet: Auxiliary Metadata Matters for Omni-Domain Infrared Small Target Detection

### v1

@@ 4.3.3 Ablation Study on Lightweight Edge Enhancement
Additionally, to probe the interplay between M2DM and LEEM, we systematically assessed their standalone and joint configurations. As evidenced in Tab. 5, their combined deployment achieves a striking +3.2% AP50 uplift, validating the architectural coherence and functional potency of our design. This synergy emerges from M2DM’s global scene-aware modulation complementing LEEM’s localized edge refinement, collectively enhancing both contextual and structural target cues.
Table 5: Ablation Study on Core Modules
Module Structure Performance Params. (M)
@@
$\checkmark$ 77.2 (+2.5) 86.7 (+2.2) 44.367 (+1.840)
$\checkmark$ 77.2 (+2.5) 86.7 (+2.2) 42.723 (+0.196)
$\checkmark$ $\checkmark$ 77.9 (+3.2) 87.2 (+2.7) 45.279 (+2.752)
4.4 Visual Analysis
Figure 3: Visualization of detection results. Green circles indicate false alarms, while green rectangles mark missed detections. Compared to other methods, AuxDet has fewer missed detections and false alarms.
As depicted in Fig. 3, we performed a qualitative comparison of five state-of-the-art methods across diverse and challenging scenarios.
  • In the first and third rows, targets are embedded within intricate backgrounds featuring highly confusable elements, such as rocks and water ripples. Competing methods, except for our AuxDet, all other methods exhibit a propensity for false alarms, misidentifying background clutter as targets.
  • In the second and fourth rows, the evaluation shifts to space-based IRSTD scenarios characterized by densely distributed small targets. Here, PConv, EFLNet, and Cascade RPN suffer from significant missed detections.
These methods, tethered to rigid visual modeling frameworks, struggle to generalize across diverse spectral bands and sensor platforms. AuxDet, however, leverages its multimodal architecture to dynamically calibrate feature representations using metadata semantics, facilitating robust domain adaptation. This approach yields substantially fewer missed detections and enhances detection fidelity across heterogeneous scenes.
In the visualization, we use red and blue masks to indicate clearly distinguishable sea areas and visually ambiguous regions with strong cloud interference in maritime detection scenes, respectively.
As observed, when the sea and land areas are relatively balanced and the sea is clearly visible to the naked eye (columns 3–4), our method provides a strong prior localization toward the sea regions. When the sea area is minimal (columns 5–6) and small land masses are visually similar to the targets, our method is still able to directly focus on the actual targets, effectively reducing false alarms. In highly entangled scenarios where sea, atmosphere, and land are difficult to distinguish (columns 7–8), our method not only maintains strong prior localization toward the sea, but also successfully highlights target regions partially obscured by clouds and barely discernible to the human eye.
5 Conclusion
In this paper, we tackle the formidable challenge of Omni-IRSTD by introducing a metadata-driven paradigm that redefines the traditional reliance on visual-only detection frameworks. Through a high-dimensional fusion module, AuxDet dynamically integrates metadata semantics with visual features, effectively addressing domain heterogeneity by tailoring representations to the unique characteristics of each input sample. Moreover, we propose the LEEM module that refines target details to enhance detection precision even amidst complex backgrounds. Extensive experiments on the WideIRSTD-Full benchmark reveal that AuxDet significantly outperforms state-of-the-art approaches, affirming the indispensable role of auxiliary metadata in improving both robustness and accuracy across diverse omni-domain IRSTD scenarios.
6 Acknowledgements
We acknowledge the Tianjin Key Laboratory of Visual Computing and Intelligent Perception (VCIP) for their essential resources. Computation is partially supported by the Supercomputing Center of Nankai University (NKSC).
Appendix A Analysis of Auxiliary Metadata
In the introduction, we delve into the critical challenge of Omni-Domain Infrared Small Target Detection (Omni-IRSTD): how to achieve a unified infrared detection paradigm across heterogeneous sensing domains. To address this challenge, we propose AuxDet, a novel auxiliary metadata-driven multimodal scene-aware optimization framework. The core idea of AuxDet lies in transcending the limitations of traditional single-modal visual analysis through metadata fusion and dynamic modulation. As shown in Tab. 6, under controlled experimental settings with strictly consistent network architecture parameters, we systematically disentangle the synergistic effects of multi-source metadata combinations on detection performance.
The experimental results demonstrate:
(1) Necessity of Metadata Integration: The inclusion of any single metadata dimension (sensor platform, image resolution, or spectral band) leads to an improvement in average precision, validating the generalizable benefits of auxiliary metadata in promoting effective domain adaptation.
(2) Optimization Principle of Metadata Combination: Intrinsically coupled metadata combinations may induce modality competition effects.
As illustrated in Fig. 1, typical cases of such coupling include: Images captured by space-based sensors are predominantly in the near-infrared (NIR) band (accounting for up to 97.93% in our training set), demonstrating a strong correlation between spectral band and platform. A large proportion of air-based platform images have a fixed resolution of 256$\times$256, leading to redundancy between resolution and platform dimensions.
Table 6: Ablation Study on Metadata Composition in AuxDet
Metadata Components Detection Performance
Sensor Resolution Band AP50 (%) $\uparrow$ Recall (%) $\uparrow$
77.2 86.7
$\checkmark$ 77.4 87.2
$\checkmark$ 77.7 87.1
$\checkmark$ 77.6 87.4
$\checkmark$ $\checkmark$ 77.7 87.0
$\checkmark$ $\checkmark$ 77.4 87.4
$\checkmark$ $\checkmark$ 77.4 86.7
$\checkmark$ $\checkmark$ $\checkmark$ 77.9 87.2
These inherent couplings between the sensor platform and other metadata types render dual-modality combinations ineffective in providing incremental discriminative guidance. For instance, in the presence of LWIR and a resolution of 256$\times$256, it remains ambiguous whether an image originates from a land-based or air-based platform. However, the inclusion of platform information resolves this ambiguity. When all three modalities are jointly introduced, platform metadata plays a key role in decoupling such strong correlations, thereby enabling precise and adaptive feature modulation strategies.
Appendix B Analysis of the Embedding Depth of Core Modules
@@
In our model design, both the Multimodal Dynamic Modulation Module (M2DM) and the Lightweight Edge Enhancement Module (LEEM) are embedded after the shallow backbone features, which are rich in detailed information. This placement is not only intuitively reasonable, but also empirically validated through comprehensive experiments. As shown in the Tab. 7, we vary the embedding depth of M2DM and LEEM while keeping other parameters and modules unchanged. The depth is counted from shallow to deep; for example, a depth of 2 indicates that the core modules are inserted after stage 1 and stage 2 of the backbone, respectively.
@@
Table 7: Ablation Study on the Embedding Depth of M2DM and LEEM Modules
@@
 M2DM  LEEM AP50 (%) $\uparrow$ Recall (%) $\uparrow$
@@
Table 8: Ablation Study on the Backbone Depth
@@
Table 9: Evaluation of General Detectors with and without AuxDet
Method Before (Baseline) After (w/ AuxDet)
Faster R-CNN (Ren et al., 2015) 31.2 33.5 34.8 (+3.6) 36.6 (+3.1)
Cascade R-CNN (Cai and Vasconcelos, 2019) 35.8 37.5 37.5 (+1.7) 39.0 (+1.5)
Dynamic R-CNN (Zhang et al., 2020a) 35.0 37.2 36.8 (+1.8) 38.8 (+2.0)
Grid R-CNN (Lu et al., 2019) 32.2 34.8 33.8 (+1.6) 36.0 (+1.2)
Libra R-CNN (Pang et al., 2019) 30.7 33.8 32.2 (+1.5) 34.8 (+1.0)
SABL (Wang et al., 2020) 64.1 77.8 65.9 (+1.8) 79.3 (+1.5)
Appendix C Analysis of Generalization and Integration Flexibility of the Proposed Framework
C.1 Analysis of the Backbone Depth
As shown in Tab. 8, our proposed AuxDet consistently outperforms the baseline across different backbone depths, demonstrating strong generalization ability and adaptability to various network capacities. Notably, the best performance is achieved when using ResNet-50 as the backbone, which strikes a balance between semantic richness and spatial detail. Shallow backbones like ResNet-18 or ResNet-34 may lack sufficient semantic abstraction, while deeper ones such as ResNet-101, despite offering richer semantics, tend to dilute fine-grained details that are critical for small target detection.
@@ C.1 Analysis of the Backbone Depth
C.2 Analysis of the Integration Flexibility of the Proposed Framework
Tab. 9 demonstrates the integration flexibility of our proposed AuxDet framework by applying it to various representative two-stage detectors. We observe consistent performance improvements in both AP50 and Recall across all detectors after incorporating AuxDet. It suggests that our method is highly compatible and complementary to existing detection pipelines.
These results validate the plug-and-play nature of AuxDet, which can be seamlessly integrated into a wide range of two-stage detection architectures. Regardless of the detector’s inherent structure or training strategy, AuxDet consistently boosts the model’s ability to localize small and dim targets, highlighting its general applicability and strong enhancement capability in infrared small object detection tasks.

### v2

@@ IV-E1 Integration with Transformer Backbones
PVTv2-B3 64.7 78.0 66.9 81.1
Swin-T 65.4 79.2 67.5 82.4
IV-E2 Comparison with Transformer-based IRSTD Methods
We also compare AuxDet with several recent Transformer-based IRSTD models, including SCTransNet, DQ-DETR, and DATransNet. As reported in Tab. IV, our method achieves superior performance in both AP50 and Recall, surpassing the best competing method (SCTransNet) by +2.4 AP50 and +20.8 Recall. This highlights the effectiveness of our metadata-guided scene-aware design in capturing both spatial and contextual cues essential for small target detection, especially in complex omni-domain scenarios.
TABLE IV: Comparison with transformer-based methods.
[…]
SCTransNet [27] 75.5 66.4
DQ-DETR [67] 70.1 81.3
DATransNet [72] 56.7 43.8
Ours 77.9 87.2
In summary, the results from both integration and comparative studies confirm that AuxDet is not only compatible with Transformer-based architectures but also capable of outperforming specialized Transformer models designed for IRSTD. This underscores the robustness and generalization ability of our framework across different model families and detection paradigms.
IV-F Ablation Study
IV-F1 Efficacy of Core Modules
To probe the interplay between M2DM and LEEM, we assessed their standalone and joint configurations. As evidenced in Tab. V, their combined deployment achieves a striking +3.2% AP50 uplift, validating the architectural coherence and functional potency. This synergy emerges from M2DM’s global scene-aware modulation complementing LEEM’s localized edge refinement, collectively enhancing both contextual and structural target cues.
TABLE V: Ablation study on core modules M2DM and LEEM. The best results are highlighted in bold.
Module Performance Params. (M)
@@ IV-F1 Efficacy of Core Modules
77.2 86.7 44.367
77.2 86.7 42.723
77.9 87.2 45.279
IV-F2 Analysis of Multi-Modal Dynamic Modulation Module
We systematically evaluate the effectiveness of dynamic modulation mechanisms by comparing different integration strategies of PSSM and PSCM. The ablation study in Table VI reveals three critical insights:
  • Independent Enhancement Pathways: PSSM’s edge-sensitive spatial warping (+1.1% AP50) compensates for target boundary erosion in thermal diffusion scenarios. PSCM’s cross-channel contextual modeling increases the performance by 1.0% (74.7% $\rightarrow$ 75.7%).
  • Parallel Fusion Superiority: Parallel fusion achieves 77.2% AP50 via complementary interaction, namely spatial modulation preserves structural integrity while channel attention suppresses clutter harmonics.
  • Parameter-Efficient Design: Shared dynamic projection layers reduce duplicate parameters, achieving 1.84M overhead (only 4.1% of baseline).
This evidence confirms our hypothesis: PSSM’s edge-aware spatial modulation and PSCM’s context-aware channel attention form orthogonal feature enhancement pathways. Their parallel implementation maximizes mutual reinforcement while avoiding parametric explosion through weight-sharing in dynamic projection layers.
TABLE VI: Ablation study on the M2DM module. The best results are highlighted in bold.
Module Config. Performance Params. (M)
AP50 (%) $\uparrow$ Recall (%) $\uparrow$
Baseline 74.7 84.5 42.527
PSSM Only 75.8 85.7 44.234
PSCM Only 75.7 85.3 44.230
PSSM $\rightarrow$ PSCM 76.3 86.1 44.367
PSCM $\rightarrow$ PSSM 76.8 86.3 44.367
PSSM + PSCM 77.2 86.7 44.367
IV-F3 Analysis of Lightweight Edge Enhancement
To substantiate the efficacy and efficiency of the LEEM, we benchmarked it against a $3\times 3$ edge convolution operator, initialized to mimic LEEM’s behavior (per Eq. (18)). As reported in Tab. VII, our decomposed 1D convolutional design delivers commensurate performance gains with a parameter increase of less than 1%.
TABLE VII: Ablation study on LEEM. The best results are highlighted in bold.
Module Config. Performance Params. (M)
AP50 (%) $\uparrow$ Recall (%) $\uparrow$
Baseline 74.7 84.5 42.527
w/ W3×3 (Standard) 77.3 86.7 43.116
w/ $_{1\times 3}\rightarrow W$3×1 77.2 86.8 42.723
IV-F4 Impact of the Embedding Depth of Core Modules
@@
In our model design, both the M2DM and the Lightweight Edge Enhancement Module (LEEM) are embedded after the shallow backbone features, which are rich in detailed information. This placement is not only intuitively reasonable, but also empirically validated through comprehensive experiments. As shown in Tab. VIII, we vary the embedding depth of M2DM and LEEM while keeping other parameters and modules unchanged. The depth is counted from shallow to deep; for example, a depth of 2 indicates that the core modules are inserted after stage 1 and stage 2 of the backbone, respectively.
@@
TABLE VIII: Ablation study on the embedding depth of M2DM and LEEM modules. The best results are highlighted in bold.
@@
M2DM-depth LEEM-depth AP50 (%) $\uparrow$ Recall (%) $\uparrow$
@@
IV-F5 Impact of the Backbone Depth
As shown in Tab. IX, our proposed AuxDet consistently outperforms the baseline across different backbone depths, demonstrating strong generalization ability and adaptability to various network capacities. Notably, the best performance is achieved when using ResNet-50 as the backbone, which strikes a balance between semantic richness and spatial detail. Shallow backbones like ResNet-18 or ResNet-34 may lack sufficient semantic abstraction, while deeper ones such as ResNet-101, despite offering richer semantics, tend to dilute fine-grained details that are critical for small target detection.
TABLE IX: Ablation study on the backbone depth. The best results are highlighted in bold.
@@ IV-F5 Impact of the Backbone Depth
IV-F6 Impact of Metadata Composition
As shown in Tab. X, under controlled experimental conditions with consistent network architecture parameters, we systematically analyze the synergistic effects of multi-source metadata integration within the Multi-Modal Dynamic Modulation Module on detection performance. The results demonstrate the necessity of metadata fusion: incorporating any single metadata dimension (sensor platform, image resolution, or spectral band) improves average precision, while their combined input further significantly enhances detection effectiveness, validating the generalizable benefits of auxiliary metadata in facilitating robust cross-domain adaptation.
TABLE X: Ablation study on metadata composition. The best results are highlighted in bold.
Metadata Components Detection Performance
[…]
74.7 84.5
76.7 86.3
76.9 86.6
76.5 86.2
77.2 86.7
V Conclusion
In this paper, we tackle the critical challenge of Omni-IRSTD by introducing a metadata-driven paradigm that redefines the traditional reliance on visual-only detection frameworks. Through a high-dimensional fusion module, AuxDet dynamically integrates metadata semantics with visual features, effectively addressing domain heterogeneity by tailoring representations to the unique characteristics of each input sample. Moreover, we propose the LEEM module that refines target details to enhance detection precision even amidst complex backgrounds. Extensive experiments on the WideIRSTD-Full benchmark reveal that AuxDet significantly outperforms state-of-the-art approaches, affirming the indispensable role of auxiliary metadata in improving both robustness and accuracy across diverse omni-domain IRSTD scenarios.
Acknowledgment
The authors would like to thank the editor and the anonymous reviewers for their critical and constructive comments and suggestions. We acknowledge the Tianjin Key Laboratory of Visual Computing and Intelligent Perception (VCIP) for their essential resources. Computation is partially supported by the Supercomputing Center of Nankai University (NKSC).